# supplement

### v1

@@ 4.1 Comparison of different gradient estimation methods
We use a simple one-dimensional problem to illustrate the effectiveness of the gradient estimation method proposed in this article. We consider a model with the prior density $\pi(\theta)=\mathcal{N}(\theta;0,1)$, the likelihood function $p(y\mid\theta)=\mathcal{N}(y;\theta,0.01)$, the observation $y_{obs}=0$, and the weighted kernel function $K_{\varepsilon}(x,y_{obs})=\mathcal{N}(x;y_{obs},\varepsilon^{2})$, where $\mathcal{N}(\theta,\mu,\sigma^{2})$ denotes a Gaussian density with mean $\mu$ and variance $\sigma^{2}$. The closed expression of estimated likelihood and gradient based on $\varepsilon$-kernel are $p_{\varepsilon}(y\mid\theta)=\mathcal{N}(y,\theta,\varepsilon^{2}+0.01)$, and $\nabla_{\theta}\log p_{\varepsilon}(y\mid\theta)=(y-\theta)/(\varepsilon^{2}+0.01)$.
@@
Figure 1: Comparison of gradient estimation methods. The shaded region represents $2\sigma$ of estimated gradient. The standard method (MC_Random), the method of Meeds et al. (2015) (MC_CRN_max, Eq. (3.12)), the improved method based on Meeds et al. (2015) (MC_CRN_mean, Eq. (3.13)), the likelihood fitting using Gaussian distribution based on simulation data with CRNs (Gaussian_CRN, Eq. (3.14) with CRNs), and the likelihood fitting using Gaussian distribution based on random simulation data (Gaussian_Random, Eq. (3.14)) are compared.
In Supplement LABEL:supp:sec-gradient, we demonstrate that for a fixed random seed $\omega_{s}$, the simulation data $y$ is deterministic for a given $\theta$ and varies smoothly with changes of $\theta$, despite the inherent noise in the simulator. For numerical methods, we fixed $d_{\theta}=0.05$. Figure 1 provides gradient estimates obtained from various methods. For each parameter value, gradients are estimated using 100 simulated data points, and the process is repeated 1,000 times to calculate both the mean and standard deviation of the gradient estimates.
Compared to the standard method (random sampling) and the $CRN_{max}$ approach (Meeds et al., 2015), the $CRN_{mean}$ method significantly reduces the variance of the estimates and provides more accurate estimates in regions with high likelihood values. However, in regions where $p_{\varepsilon}(y\mid\theta)$ approaches zero, substantial deviations are observed because the logarithmic function amplifies the errors in the Monte Carlo estimates. By employing the CRNs technique, fitting a Gaussian density can significantly enhance both the accuracy and variance of the estimates.
@@ 4.2 Three synthetic probability densities
We illustrate the advantages of our approach through three synthetic examples: Mixture Gaussian, Moon and Wave shaped posterior. The visualization of these posterior densities are shown in Figure 2. In this simulation, MCMC with a local proposal involves generating candidate particle $\theta^{*}$ from a normal distribution $N(\theta_{t-1},\sigma^{2})$ at the $t$-th iteration. On the other hand, MCMC with a global proposal (MCMCg) generates candidate particle $\theta^{*}$ from a proposal distribution $q(\cdot)$ that is independent with $\theta_{t-1}$, at the $t$-th iteration. The true ABC posterior is estimated using $5\times 10^{5}$ samples, obtained by running importance sampling with $10^{8}$ samples from the prior. We use the bkde2D function from the R package KernelSmooth (Wand, 1994) to estimate the posterior density based on the posterior samples. The accuracy of different methods is measured by the KL divergence between the estimated posterior and the true posterior, calculated as
where $\pi_{\varepsilon}$ is the true ABC posterior density and $\hat{\pi}$ is the estimated ABC posterior density. In Figure 2, $\Theta_{ref}$ is generated in a two-dimensional grid of $501^{2}$ points on the posterior region, that is $[-3,3]\times[-3,3]$ of Mixture, $[-2,2]\times[-5,1]$ of Moon, and $[-1,1]\times[-4,4]$ of Wave. More details of each simulation model are provided in Supplement LABEL:supp:sec-toy.
@@
Firstly, we utilize the Mixture example to illustrate how ABC-i-SIR can improve the mixing of ABC-MCMC with local proposals, which often fails to switch effectively between peaks. Figure 3 displays a segment of the trace plots (i.e., iteration $30,001\sim 40,000$) for three MCMC methods: ABC-MCMC with a Gaussian proposal (left), ABC-MCMC with prior distribution as the global proposal (middle), and ABC-i-SIR with prior distribution as the importance proposal, with batch size $51$ (right). The analysis reveals that ABC-MCMC with local proposals struggles to transition efficiently between peaks, while ABC-MCMC samplers with global proposals can traverse between peaks, their efficiency in exploring the local posterior region is notably low. In contrast, the ABC-i-SIR algorithm demonstrates that its particles are capable of both traversing between peaks and effectively exploring the posterior region.
@@
Additionally, we employ the Mixture example to investigate the influence of various proposal distributions and batch size on the performance of ABC-i-SIR. We use three different distributions as the global proposal of MCMC and the importance proposal of ABC-i-SIR respectively. There are, Uniform: ${U}(-4,4)^{2}$, Prior: $N(0,I_{2})$ and Optimal: $(N((1.425,1.425)^{\intercal},0.28^{2}I)+N((-1.425,1.425)^{\intercal},0.28^{2}I)+N((1.425,-1.425)^{\intercal},0.28^{2}I)+N((-1.425,-1.425)^{\intercal},0.28^{2}I))/4$.
@@
Figure 4 shows that the accuracy of the ABC-i-SIR algorithm increases with the batch size. The selection of proposal distributions impacts the convergence speed of the ABC-i-SIR algorithm, but this effect diminishes as the batch size increases. This indicates that we can improve the performance of the algorithm by increasing the batch size when it is difficult to construct a good proposal distribution, which could be highly advantageous for parallel simulation models. Notably, ABC-MCMC with global proposals outperforms ABC-i-SIR with batch size 1.
In cases where generating simulation data is constrained, we can collect training data during the algorithm’s iterative process and employ normalizing flows to enhance the proposal distribution. Figure 5 demonstrates that NFs can significantly improve the algorithm’s convergence speed. Here the proposal is the prior distribution, and we update the NF model every time we collect 1000 simulated data points, with a total of 50 updates.
Figure 5: Comparison of different methods on whether to perform NFs training. ‘$\text{MCMC}_{g}$’: MCMC with a global proposal. ‘$i$’: iSIR with batch size $i$. Dashed line: methods with NF, solid line: methods without NFs.
Tables 1 presents a comparison of several ABC-MCMC methods for the Mixture, Moon, and Wave examples. GL-ABC-MCMCs are the combinations of MALA and ABC-i-SIR, using the prior distribution as the importance proposal. The step size of MALA is chosen using a sequential optimal algorithm based on ESJD. Notably, GL-ABC-MCMC degenerates into MALA when $\gamma=0$ and into ABC-i-SIR when $\gamma=1.0$. Each MCMC algorithm was repeated10 times, with 1,000,000 iterations per run. Table 1 shows the KL divergences between the estimated posterior of different ABC-MCMC methods and the ground truth, along with their standard deviations. In the three examples, the combination of global proposal and local proposal demonstrated superior performance compared to ordinary ABC-MCMC and MALA algorithms, especially in the case of the multi-modal Mixture example. Compared to the standard MCMC algorithm, the MALA algorithm, which utilizes the gradient of the log-likelihood function, achieved a more accurate posterior estimate and a larger effective sample size. This was especially true for models with posterior distributions exhibiting strong correlations, such as Moon and Wave examples (detailed results can be seen in Supplement LABEL:supp:sec-toy, Table LABEL:tab:Toy_ESS). When using a batch size of 5, the GL-ABC-MCMC algorithm with $\gamma=0.4$ in the Mixture example and $\gamma=0.8$ in the Moon example performed the best. This suggests that the combination of global and local proposals may be particularly effective when the global exploration ability of the algorithm is relatively weak.
@@
Table 1: The KL divergence between the estimated posterior of different ABC-MCMC methods and the ground truth for Moon and Banana data sets. $\gamma$ denotes global frequency and ‘bsi’ denotes batch size $i$.
Method Mixture $({10^{-3}(10^{-4})})$ Moon ($10^{-3}(10^{-4})$) Wave ($10^{-5}(10^{-6})$)
@@
MCMC 1658.5(1145.9) 13.1(85.4) 63.1(470.7)
MALA 1650.0(922.0) 5.9 (33.3) 18.3(86.6)
@@ 5.1 Example 1: Roman binary microlensing events
A binary-lens, single-source (2L1S) microlensing event is an astronomical phenomenon where the light from a distant source star is bent and magnified as it passes through a binary star system. This phenomenon offers a unique opportunity to study and discover exoplanets. Fast and automated inference of 2L1S microlensing events using MCMC methods faces two main challenges: (a) the high computational cost of likelihood evaluations using microlensing simulation codes, and (b) the complexity parameter space characterized by a negative-log-likelihood surface riddled with numerous narrow and deep local minima.
This section examines the 2L1S microlensing model explored in Zhang et al. (2021). In our analysis, we fix the angle of the source trajectory relative to the projected binary lens axis at $\alpha=110$ degrees and the time of closest approach at $t_{0}=60.0$ day. The model comprises five unknown parameters in the model: binary lens separation ($s$), mass ratio ($q$), impact parameter ($u_{0}$), Einstein ring crossing timescale ($t_{E}$), and source flux fraction ($f_{s}$).
We simulate the 2L1S magnification sequences using the microlensing code MulensModel (Poleski and Yee, 2019) over a 144 days period with a cadence of 15 minutes, resulting in $N_{\boldsymbol{y}}=13825$ observations. The observed data $\boldsymbol{y}$ is generated using the parameters $(s,q,u_{0},t_{E},f_{s})=(10^{-0.2},10^{-2.5},0.2,10^{1.6},0.2)$. The discrepancy between the observed data $\boldsymbol{y}$ and the simulated data $\boldsymbol{x}$ is defined as $\Delta(\boldsymbol{x},\boldsymbol{y})=\frac{1}{N_{\boldsymbol{y}}}\sum_{i=1}^{N_{\boldsymbol{y}}}|\boldsymbol{y}_{i}-\boldsymbol{x}_{i}|$. A Gaussian kernel with a bandwidth $\varepsilon=0.003$ is employed. We simulate 2L1S events using the analytical priors: $s\sim\text{LogUniform}(0.2,0.5)$, $q\sim\text{LogUniform}(10^{-6},1)$, $u_{0}\sim\text{Uniform}(0,2)$, $t_{E}\sim\text{TruncLogNorm}(1,100,\mu=10^{1.15},\sigma=10^{0.45})$, and $f_{s}\sim\text{LogUniform}(0.1,1)$.
For local proposals, we use the following distributions: $\log(s^{*})\sim N(\log(s),0.2^{2})$, $\log(q^{*})\sim N(\log(q),0.2^{2})$, $u_{0}^{*}\sim N(u_{0},0.01^{2})$, $\log(t_{E}^{*})\sim N(t_{E},0.05^{2})$, and $\log(f_{s}^{*})\sim N(f_{s},0.05^{2})$. Due to the model’s complexity, we derive the importance proposal from a coarse posterior estimate obtained by running 5,000 iterations of simple rejection sampling with a large threshold of $\varepsilon=0.05$.
Considering simulation costs, a sequential optimization algorithm is employed to determine the global frequency $\gamma$ and the batch size $N_{b}$, constrained by $\gamma N_{b}+(1-\gamma)=C$, to maintain the average number of simulation per iteration to $C=3$. For each potential set of hyperparameters, we perform GL-ABC-MCMC with $5,000$ iterations, repeating this process five times.
@@
Figure 6 presents the posterior estimates obtained via basic ABC-MCMC with a random walk proposal, ABC-MCMC with the importance proposal as the global proposal (MCMCg), GL-ABC-MCMC, ABC-iSIR with batch size 3, independent multiple try Metropolis (IMTM) (Martino, 2018) and the likelihood-free approximate Gibbs (Rodrigues et al., 2020). The basic ABC-MCMC, MCMCg and GL-ABC-MCMC are all executed over 500,000 iterations. We use a Gaussian process model to approximate the full conditional distributions, and $50,000$ train data is from importance proposal. We run the approximate Gibbs 110,000 iterations and discard the first 10,000 iterations as burn-in. Firstly, ABC-MCMC with global proposal, ABC-iSIR and independent multiple try Metropolis (IMTM) with batch-size 3 present random walk behavior, and the approximate Gibbs fails to capture the multiple modes of the posterior distribution. Both the standard ABC-MCMC and GL-ABC-MCMC can effectively explore both peaks of the posterior. Subsequently, we compare the effective sample size (ESS) of GL-ABC-MCMC against that of traditional ABC-MCMC approaches. ABC-MCMC1 is configured with the same number of iterations as GL-ABC-MCMC, while ABC-MCMC2 is adjusted to have a comparable number of simulations as GL-ABC-MCMC. According to Table 2, GL-ABC-MCMC demonstrates a significantly larger ESS compared to both ABC-MCMC1 and ABC-MCMC2, indicating superior performance of our GL-ABC-MCMC approach. Table 3 shows the estimates of the posterior mean and credible intervals for different methods. The posterior mean obtained from GL-ABC-MCMC is closer to the true value, and the GL-ABC-MCMC estimate exhibits a narrower 95% credible interval.
Table 2: Comparison of effective sample sizes of different methods. ABC-MCMC1: $N=500,000$; ABC-MCMC2: $N=1,500,000$; GL-ABC-MCMC: $N=500,000$ and the average number of simulations per iteration is $3$.
@@
ABC-MCMC1 579.84 75.10 356.57 405.17 382.94
ABC-MCMC2 1768.85 229.93 973.57 1096.42 1019.30
GL-ABC-MCMC 1389.22 748.04 2235.37 2270.99 2172.30
Table 3: Comparison of posterior mean and 95% credible intervals of different methods.
Method $\log(s)$ $\log(q)$ $u_{0}$ $log(t_{E})$ $log(f_{s})$
Ture -0.4605 -5.7565 0.2 3.6841 -1.6094
ABC-MCMC1 -0.1716 -8.4197 0.2215 3.6402 -1.5183
(-1.55,1.49) (-13.49,-2.39) (0.15,0.33) (3.37,3.90) (-1.94,-1.01)
ABC-MCMC2 -0.1079 -8.4456 0.2240 3.6349 -1.5082
(-1.55,1.51) (-13.51,-2.79) (0.15,0.34) (3.36,3.91) (-1.97,-0.99)
GL-ABC-MCMC -0.1657 -8.3515 0.2158 3.6577 -1.5500
(-1.56,1.50) (-13.57,-1.76) (0.14,0.32) (3.39,3.93 ) (-2.00,-1.07)

### v2

@@ 5.1 Comparison of different gradient estimation methods
We use a simple one-dimensional problem to illustrate the effectiveness of the gradient estimation method proposed in this article. We consider a model with the prior density $\pi(\theta)=\mathcal{N}(\theta;0,1)$, the likelihood function $p(y\mid\theta)=\mathcal{N}(y;\theta,0.01)$, the observation $y_{obs}=0$ and the weighted kernel function $K_{\varepsilon}(x,y_{obs})=\mathcal{N}(x;y_{obs},\varepsilon^{2})$, where $\mathcal{N}(\theta,\mu,\sigma^{2})$ denotes a Gaussian density with mean $\mu$ and variance $\sigma^{2}$. Then the closed expression of estimated likelihood and gradient based on $\varepsilon$-kernel are $p_{\varepsilon}(y\mid\theta)=\mathcal{N}(y,\theta,\varepsilon^{2}+0.01)$, and $\nabla_{\theta}\log p_{\varepsilon}(y\mid\theta)=(y-\theta)/(\varepsilon^{2}+0.01)$.
@@
Figure 1: Comparison of gradient estimation methods. The shaded curved region represents $2\sigma$ of estimated gradient. The standard method (MC_Random), the method of meeds2015HABC (MC_CRN_max, Eq. (4)), the improved method based on meeds2015HABC (MC_CRN_mean, Eq. (5)), the likelihood fitting using Gaussian distribution based on simulation data with CRNs (Gaussian_CRN, Eq. (6) with CRNs), and the likelihood fitting using Gaussian distribution based on random simulation data (Gaussian_Random, Eq. (6)) are compared.
In the Appendix C.1, we demonstrate that for a fixed random seed $\omega_{s}$, the simulation data $y$ is deterministic for a given $\theta$ and varies smoothly with changes of $\theta$, despite the inherent noise in the simulator. For numerical methods, we fixed $d_{\theta}=0.05$. Figure 1 provides gradient estimates obtained from various methods. For each parameter value, gradients are estimated using 100 simulated data points, and repeat the process 1,000 times to calculate both the mean and standard deviation of the gradient estimates.
Compared to the standard method (random sampling) and the $CRN_{max}$ approach (meeds2015HABC), the $CRN_{mean}$ method significantly reduces the variance of the estimates and provides more accurate estimates in regions with high likelihood values. However, in regions where $p_{\varepsilon}(y\mid\theta)$ approach zero, substantial deviations are observed because the logarithmic function amplifies the errors in the Monte Carlo estimates. By employing the CRNs technique, fitting a Gaussian density can significantly enhance both the accuracy and variance of the estimates.
@@ 5.2 Three synthetic probability densities
We illustrate the advantages of our approach through three synthetic examples, such as Mixture Gaussian, Moon and Wave shaped posterior. The visualization of posterior densities are shown in Figure 2. In this simulation, MCMC with a local proposal means the candidate particle $\theta^{*}\sim N(\theta_{t-1},\sigma^{2})$ at the $t$-th iteration, and MCMC with global proposal (MCMCg) means the candidate particle $\theta^{*}\sim q(\cdot)$, where $q(\cdot)$ is independent with $\theta_{t-1}$, at the $t$-th iteration. The true ABC posterior is estimated using $5\times 10^{5}$ samples, which is obtained by running importance sampling with $10^{8}$ samples from the prior. We use the bkde2D function in the R package KernelSmooth (wand1994fast; wand1994kernel) to estimate the posterior density based on the posterior samples. The accuracy of different methods is measured by the KL divergence between the estimated posterior and the true posterior, which is calculated as
where $\pi_{\varepsilon}$ is the true ABC posterior density and $\hat{\pi}$ is the estimated ABC posterior density. In Figure 2, $\Theta_{ref}$ is generated in a two-dimensional grid of $501^{2}$ points on the posterior region, that is $3,3]\times[-3,3]$ of Mixture, $2,2]\times[-5,1]$ of Moon, and $1,1]\times[-4,4]$ of Wave. More details of each simulation model are provided in the Appendix C.2 of the Supplementary Material.
@@
Firstly, we utilize the mixture Gaussian example to illustrate how ABC-i-SIR can improve the mixing of ABC-MCMC with local proposals, which fails to switch effectively between peaks. Figure 3 displays part of the trace plots (i.e., iteration $30,001\sim 40,000$) for three MCMC methods: ABC-MCMC with a Gaussian proposal (left), ABC-MCMC with prior distribution as the global proposal (middle), and ABC-i-SIR with prior distribution as the importance proposal with batch size $51$ (right). The analysis reveals that ABC-MCMC with local proposals struggle to transit efficiently between peaks, while ABC-MCMC samplers with global proposals can traverse between peaks, their efficiency in exploring the local posterior region is notably low. In contrast, the ABC-i-SIR algorithm demonstrates that its particles are capable of both traversing between peaks and effectively exploring the posterior region.
Additionally, we employ the mixture Gaussian example to investigate the influence of various proposal distributions and batch size on the performance of ABC-i-SIR. Here, we use three different distributions as the global proposal of MCMC and the importance proposal of ABC-i-SIR respectively, that is, Uniform: ${U}(-4,4)^{2}$, Prior: $N(0,I_{2})$ and Optimal: $(N((1.425,1.425)^{\intercal},0.28^{2}I)+N((-1.425,1.425)^{\intercal},0.28^{2}I)+N((1.425,-1.425)^{\intercal},0.28^{2}I)+N((-1.425,-1.425)^{\intercal},0.28^{2}I))/4$.
@@
Figure 4 shows that the accuracy of the ABC-i-SIR algorithm increases with the batch size. The selection of proposal distributions impacts the convergence speed of the ABC-i-SIR algorithm, but this effect gradually decreases as the batch size increases. This indicates that we can improve the performance of the algorithm by increasing the batch size when it is difficult to construct a good proposal distribution. This could be highly advantageous for parallel simulation models. It is also worth noting that ABC-MCMC with global proposals outperforms ABC-i-SIR with batch size 1.
Figure 5 presents a comparison of inference results for the Mixture Gaussian model, obtained using the Global kernel with and without adaptive updates of the importance proposal distribution. The performance of ABC-i-SIR with various batch sizes is also evaluated. Two strategies for adaptively updating the importance proposal distribution are considered: the ABC-Ai-SIR algorithm proposed in this work, and an ABC extension of samsonov2022local-global (see Appendix Section A.1 for details). The initial proposal distribution is set as the prior, and the importance proposal distribution is updated by collecting every $2,000$ simulated data points. In the ABC-Ai-SIR algorithm, the auxiliary threshold reduction factor is set to $\alpha=0.8$, the target auxiliary threshold is $\varepsilon^{*}_{T}=5\varepsilon$, and the normalizing flow (NF) model is updated 50 times.
Figure 5: Comparison of different methods regarding the use of global proposal updates for the Mixture Gaussian model. ‘bsii’ denotes ABC-i-SIR with batch size ii. Green dashed lines: methods employing NF; red solid line: the proposed ABC-Ai-SIR algorithm; blue long-dashed lines: methods without global proposal updates.
As shown in Figure 5, effective proposal updates can significantly improve inference results, and our proposed method demonstrates superior efficiency. Our proposed method only takes three to four minutes, whereas the NF method takes approximately half an hour. All experiments were conducted on a Mac mini equipped with an Apple M4 processor and 16 GB of RAM. Figure 6 illustrates the evolution of the proposal distributions of ABC-Ai-SIR. It can be observed that the proposal distribution is steadily updated towards the target distribution over iterations. Furthermore, the NF method directly uses the ABC posterior distribution as the update target. Due to the significant difference between the initial and target distributions, insufficient training data may lead to limited improvement in the proposed distribution, thus affecting inference performance. Further discussions on this issue, encompassing the evolution of proposal constructed by NF and detailed results concerning the Moon and Wave models, are provided in the Appendix C.3.
Figure 6: Evolution of the importance proposal distribution in the ABC-Ai-SIR algorithm at the 1st, 6th, 11th, and 16th updates, shown from left to right.
@@
Table 1: The KL divergence between the estimated posterior of different ABC-MCMC methods and the ground truth for Moon and Banana data sets. $\gamma$ denotes global frequency and ‘bsi’ denotes batch size ii.
Mixture $({10^{-3}(10^{-4})})$ Moon ($10^{-3}(10^{-4})$) Wave ($10^{-5}(10^{-6})$)
MCMC 1658.5(1145.9) 13.1(85.4) 63.1(470.7)
MALA 1650.0(922.0) 5.9 (33.3) 18.3(86.6)
@@
Tables 1 presents a comparison of several ABC-MCMC methods for the Mixture Gaussian Moon, and Wave. GL-MCMCs are the combinations of MALA and ABC-i-SIR with a prior distribution as the importance proposal. The step size of MALA is chosen by the sequential optimal algorithm based on ESJD. Notably, GL-MCMC degenerates into MALA when $\gamma=0$ and ABC-i-SIR when $\gamma=1.0$. Here, we repeated the MCMC algorithm 10 times, each with 1,000,000 iterations. Table 1 shows the KL divergences between the estimated posterior of different ABC-MCMC methods and the ground truth and their standard deviations. In the three examples mentioned, the combination of global proposal and local proposal has shown superior performance compared to ordinary ABC-MCMC and MALA algorithms, especially in the case of the multi-modal Mixture Gaussian example. Compared with the MCMC algorithm, the MALA algorithm, which utilizes the gradient of the log-likelihood function, has a more accurate posterior estimate and a larger effective sample size (i.e. Table 2), especially for models where the posterior distribution exhibits strong correlations, such as Moon and Wave. When using a batch size of 5, the GL-MCMC algorithm with $\gamma=0.4$ in the Mixture Gaussian example and $\gamma=0.8$ in the Moon example performs the best. This suggests that the combination of global and local proposals may be particularly effective when the global exploration ability of the algorithm is relatively weak.
Table 2: Effective sample size of the posterior samples obtained by MCMC and MALA.
Mixture Moon Wave
$\theta_{1}$ $\theta_{2}$ $\theta_{1}$ $\theta_{2}$ $\theta_{1}$ $\theta_{2}$
MCMC 4050 3940 150 164 451 956
MALA 7589 7788 962 691 1474 2710
@@ 6.1 Example 1: Roman binary microlensing events
A binary-lens, single-source (2L1S) microlensing event is an astronomical phenomenon where the light from a distant source star is bent and magnified as it passes through a binary star system. This phenomenon offers a unique opportunity to study and discover exoplanets. Fast and automated inference of 2L1S microlensing events using Markov Chain Monte Carlo (MCMC) methods faces two main challenges: (a). the high computational cost associated with likelihood evaluations using microlensing simulation codes, and (b). a complex parameter space characterized by a negative-log-likelihood surface with numerous narrow and deep local minima.
In this section, we study the 2L1S microlensing model investigated in zhang2021real. We fix the angle of the source trajectory relative to the projected binary lens axis at $\alpha=110$ degrees and the time of closest approach at $t_{0}=60.0$ day. There are in total five unknown parameters in the model: binary lens separation ($s$), mass ratio ($q$), impact parameter ($u_{0}$), Einstein ring crossing timescale ($t_{E}$), and source flux fraction ($f_{s}$).
The 2L1S magnification sequences are simulated using the microlensing code MulensModel (poleski2019modeling), over a span of 144 days with a cadence of 15 minutes, the number of observations $N_{\boldsymbol{y}}=13825$. The observed data $\boldsymbol{y}$ is generated using the parameters $(s,q,u_{0},t_{E},f_{s})=(10^{-0.2},10^{-2.5},0.2,10^{1.6},0.2)$. The discrepancy between the observed data $\boldsymbol{y}$ and the simulated data $\boldsymbol{x}$ is defined as $\Delta(\boldsymbol{x},\boldsymbol{y})=\frac{1}{N_{\boldsymbol{y}}}\sum_{i=1}^{N_{\boldsymbol{y}}}|\boldsymbol{y}_{i}-\boldsymbol{x}_{i}|$. The kernel function employed is a Gaussian kernel with a bandwidth $\varepsilon=0.003$. We then simulate 2L1S events using the analytical priors:
[…]
In all schemes that employ a local proposal, the local proposal is set to
[…]
Figure 7: Sequential optimization of hyper-parameter $N_{b}$ with constrained condition $C=5$.
First, we select the hyperparameters for our GL-ABC-MCMC method. We employ ABC-i-SIR with the prior distribution (i.e., Eq. (13)) as the importance proposal. To manage the simulation cost, we fix the average number of simulations per iteration to $C=5$, which constrains the global frequency $\gamma$ and the batch size $N_{b}$ via $\gamma N_{b}+(1-\gamma)=C$. For each candidate hyperparameter configuration, GL-ABC-MCMC is run for 5,000 iterations, and the entire procedure is repeated five times to assess variability. Figure 7 illustrates the selection of $N_{b}$ through sequential optimization, which leads us to set $N_{b}=70$. From this point onward, for all GL-ABC-MCMC runs in this example, $N_{b}$ is fixed at 70 unless otherwise noted.
@@
Figure 8 presents the posterior estimates obtained via six MCMC methods with over 500,000 iterations:
  • MCMC: Basic ABC-MCMC utilizing the proposal defined in Eq. (14).
  • MCMCg: ABC-MCMC with the proposal given by Eq. (15).
  • ABC-i-SIR: Importance sampling with $N_{b}=5$ and proposal distribution specified by Eq. (15).
  • IMIM: Independent Multiple Try Metropolis (MARTINO2018134), employing the same sets as ABC-i-SIR.
  • GL-ABC-MCMC: Global-local ABC-MCMC where the global move corresponds to ABC-i-SIR with $N_{b}=70$, using Eq. (13) as the importance proposal distribution. The probability of a global move, $\gamma$, satisfies $\gamma N_{b}+(1-\gamma)=5$.
  • Adaptive GL-ABC-MCMC: Extends GL-ABC-MCMC by adaptively updating the proposal according to Algorithm 3, with auxiliary target threshold $\varepsilon_{T}^{*}=0.015$, threshold decay parameter $\alpha=0.6$, and the importance proposal updated every $S=50$ global proposals.
The substantial discrepancy between prior and posterior leads to negligible acceptance rates when the prior is used as the global proposal, as observed in MCMCg, IMIM, and ABC-i-SIR. To address this, we construct an efficient proposal distribution
$q({\boldsymbol{\theta}})=\mathcal{N}({\boldsymbol{\theta}};\boldsymbol{\mu},\Sigma),$ (15)
where $\mathcal{N}({\boldsymbol{\theta}};\boldsymbol{\mu},\Sigma)$ denotes a Gaussian distribution with mean vector $\boldsymbol{\mu}$ and covariance matrix $\Sigma$. The parameters $\boldsymbol{\mu}$ and $\Sigma$ are estimated from the 500 realizations with the smallest discrepancies among 5,000 prior draws.
As illustrated in Figure 8, ABC-MCMC with a global proposal, ABC-i-SIR, and IMIM exhibit random-walk behavior, hindering efficient exploration. In contrast, the standard ABC-MCMC, GL-ABC-MCMC, and adaptive GL-ABC-MCMC successfully explore both modes of the posterior distribution.
Table 3: Comparison of effective sample sizes of different methods: ABC-MCMC, GL-ABC-MCMC, GL-ABC-MCMCc (using Eq. (15) as the importance proposal), Adptive GL-ABC-MCMC. The average number of simulations per iteration of GL-ABC-MCMC is $5$.
@@
ABC-MCMC 1052.428 132.8922 206.6277 271.4922 228.3610
GL-ABC-MCMC 1133.471 314.4186 481.5078 476.1921 464.2407
GL-ABC-MCMCc 2252.269 1183.1576 1863.4926 1649.8262 1647.7599
Aaptive GL-ABC-MCMC 7516.687 6630.4125 8273.9112 7362.3383 7501.7451
Figure 9: Posterior mean trajectories and 95% credible intervals versus iteration number for various ABC-MCMC methods. The black dotdash lines are true values we used to simulate the data.
Table 3 compares the effective sample sizes (ESS) across the standard ABC-MCMC, GL-ABC-MCMC, GL-ABC-MCMCc and adaptive GL-ABC-MCMC approaches. Figure 9 examines the evolution of the estimated posterior means as a function of MCMC iterations. The GL-ABC-MCMCc is constructed with importance proposal shown in Eq. (15). The results show that GL-ABC-MCMC strategies outperform standard ABC-MCMC that relies on local proposals in terms of both ESS and convergence speed. For this complex model, a global move constructed from the prior offers limited improvement and falls short of the performance achieved by the constructed proposal distribution. The adaptive method we proposed exhibits substantially superior performance: with the same number of iterations, the ESS is tens of times larger than that of ABC-MCMC, and the posterior means stabilize rapidly.
We attempted to use normalizing flows to update the proposal distribution under the same parameter settings as the GL-ABC-MCMC algorithm. Since the posterior region is much narrower compared to the prior, only about one hundred samples with non-zero weights were obtained after $50$ global moves (i.e., $3500$ samples in total). The limited number of effective training samples caused issues in training the normalizing flow, resulting in the support of the proposal distribution deviating from the posterior region. Consequently, the global moves could not perform effectively. Therefore, the results of NF method is not used for comparison. Nevertheless, in the Appendix Figure C.4, we present the evolution of the proposal distributions using both the method proposed in this work and the NF method. During the training of the NF, if 50 global training iterations pass without obtaining 500 training samples with importance weights greater than zero, the training is paused until 500 such samples are collected. The importance proposals constructed by the NF method fail to fully cover the posterior region, further highlighting the necessity and effectiveness of introducing intermediate target distributions when dealing with complex posterior models. More details of this experiment are described in Appendix C.4.